Find 3 consecutive numbers such that 12 less than three times the first number is the same as the sum of the other 2 numbers.
Answer by mananth(16949)

let x, x+1, x+2 be the integers
3x-12= x+1+x+2
3x-12=2x+3
3x-2x=3+12
x=15
15,16,17